the slope of a line is 1/2. If the line passes through the points (0, k) and (5, 4), what is the value of k?

Answer by stanbon(75887)

the slope of a line is 1/2. If the line passes through the points (0, k) and (5, 4), what is the value of k?
---------
Equation:
(k-4)/(0-5) = 1/2
---------
k-4 = -5(1/2)
------
k-4 = -5/2
---
k = -5/2 + 8/2
-----
k = 3/2
===================
Cheers,
Stan H.